OpenStudy (oswaldmurphy):

Choose the correct simplification of the expression −3x2(5x − 4x2 − 6). −7x4 + 2x3 − 9x2 12x4 − 15x3 + 18x2 12x4 + 15x3 + 18x2 −12x4 + 15x3 − 18x2 Can someone help me understand this?

OpenStudy (oswaldmurphy):

@phi

OpenStudy (phi):

you should re-order this part (5x − 4x^2 − 6) so the largest exponent is first: (-4x^2 + 5x - 6) now multiply each term by −3x^2 −3x^2 (-4x^2 + 5x - 6) do you know what -3x^2 * -4x^2 is ?

OpenStudy (phi):

remember -3x^2 * -4x^2 means -3*-4*x^2 *x^2 what is -3 * -4 ?

OpenStudy (oswaldmurphy):

12

OpenStudy (oswaldmurphy):

12^2 or 12^4?

OpenStudy (phi):

-3*-4 is 12 so -3*-4*x^2*x^2 is the same as 12*x^2*x^2 remember x^2 means x*x so we could write 12*x^2*x^2 is the same as 12*x*x*x*x

OpenStudy (phi):

most people use the "short-cut" way to write x*x*x*x (using a "little 4" in the upper right of x)

OpenStudy (oswaldmurphy):

Ok. Do you add or multiply the exponents?

OpenStudy (phi):

it's easier to remember, if you get the idea that the "exponent" way is just a short cut. for example: x^1 * x^2 means x * x * x and x*x*x is x^3 using the "short-cut" notice the exponent 3 is 1+2

OpenStudy (phi):

if you can't remember any of that, and you see x^2 *x^2 write it as x*x * x*x now count the x's and use the short-cut way to write

OpenStudy (oswaldmurphy):

Ok, but my teacher has to talk to me and ask questions. 1 might be, if you multiply exponents do you add them or multiply them? I just wanna know for the future

OpenStudy (phi):

yes. And I am saying you will remember the answer better if it "makes sense" and it will make sense if you can figure it out. And I know you can figure it out. Just go slow with this problem: x^1 * x^2 = x * x*x = x^3 notice the 3 is the sum of 1+2

OpenStudy (phi):

−3x^2 (-4x^2 + 5x - 6) the first term will be 12x^4 now what do you get for the next term: -3x^2 * 5x ? any idea ?

OpenStudy (oswaldmurphy):

15x^2?

OpenStudy (phi):

what is -3*5 ?

OpenStudy (oswaldmurphy):

-15 sorry XD -15x^2

OpenStudy (phi):

-3x^2 * 5x is -3*x*x * 5*x or re-arranging it -3*5*x*x*x now simplify -3*5 and write x*x*x using exponents

OpenStudy (oswaldmurphy):

x^3

OpenStudy (phi):

yes, that means -3x^2 * 5x = -15x^3

OpenStudy (phi):

the last term is −3x^2 * -6

OpenStudy (phi):

−3x^2 * -6 re-arrange it as -3*-6*x*x

OpenStudy (oswaldmurphy):

-18x^2

OpenStudy (phi):

you have trouble with minus signs: if you multiply minus times minus you get plus

OpenStudy (oswaldmurphy):

Wait 18x^2 I just noticed

OpenStudy (oswaldmurphy):

I didnt see the 6 was -.. sorry :P

OpenStudy (oswaldmurphy):

so the answer is B?

OpenStudy (phi):

ok, now you have the answer −3x^2 (-4x^2 + 5x - 6) = 12x^4 -15x^3 +18x^2

OpenStudy (oswaldmurphy):

Thank you :D